### 30 Jan 2024 v1

@@ 6.2.1 The edge of the HII region
In our work, we estimate the POS magnetic field strength at the border of the expanding HII region to be µG on the west side, and µG on the east side. The good agreement between the CN and HCO strengths in these regions, particularly at the eastern side, could indicate that these two molecular lines are tracing similar gas. Nevertheless, the factor of almost 2 difference in the measured in the west and the east could be due to different gas densities at these sides of the bubble. As the eastern side of the bubble is denser than the western side (see e.g. Meyer et al., 2008), we expect that the gas here is less impacted by the incoming radiation. In addition, the opacity of the HCO emission is notably higher () on the western side than on the eastern side ().
@@ 6.2.2 Filament
Overall, our measurement of the magnetic field strength in the filament is lower than in other filamentary structures. For example, Pattle et al. (2017) measured a POS magnetic field strength of a few mG in the Orion A filament. Similarly, as discussed in the previous section, it is important to highlight that Pattle et al. (2017) used CO emission to measure the turbulence and number density and different method to derive the . Similarly, Ching et al. (2018) investigated magnetic field towards the DRL1 filament and found µG.
6.3 Is the magnetic field sufficient to prevent star formation in NGC 2024?
@@
We report our results in Tab. 4 for values of magnetic field strengths computed from CN and HCO respectively using approach from Skalidis & Tassis (2021); Skalidis et al. (2021). We estimate in each region. This result could imply the presence of magnetically supported gas (see e.g. Pattle et al., 2019). In general, it is worth noting that we see a difference of one and two orders of magnitude in the between the edges of the bubble and the filament and the overlap region.
@@
The results reported in Tab. 4 should be taken as upper limits considering we calculated them using the POS component of the magnetic field. In particular, the reported mass-to-flux ratio must be taken with caution. The complex geometry of NGC 2024, large uncertainties of reported POS magnetic field strength, and non availability of the LOS component of the -field have a big impact on values reported in Tab. 4. Our results imply that the edges of the bubble show different properties than the filamentary structure and the overlap region in NGC 2024. These regions are indeed different in terms of the impact of stellar feedback. However, further quantification of gravitational stability of the gas in these regions require systematic analysis of the magnetic field. This also includes both POS and LOS components in these regions and constraints on the geometry of NGC 2024, which is beyond the scope of this study.
The properties of the gas we study in our work vary with the environment and this gas shows different characteristics. Our results thus highlight the non-negligible role of the magnetic field in NGC 2024 in regions impacted by the stellar feedback. The edges of the bubble in NGC 2024 are magnetically-dominated and gas in these regions is magnetically supported against gravitational collapse. However, the high values of the Alfvénic Mach number suggest that these regions are highly turbulent. The gravitational stability of gas located at the shell of expanding HII region is not always the case. For example, the expansion of the HII region can also trigger star formation, as seen in Galactic HII regions RCW 82 (Pomarès et al., 2009) and RCW 120 (Figueira et al., 2017). Since the HII region in NGC 2024 is relatively young, it is possible that the stellar feedback has not yet been strong enough to trigger star formation. Nevertheless, it is crucial to point out that our result is based on CN and HCO measurements, which trace the UV-illuminated and UV-shielded gas, which do not necessarily trace the star-forming gas.
@@ 7 Conclusions
We present new SOFIA HAWC+ dust polarization measurements across the NGC 2024 HII region and associated molecular cloud in the Orion B molecular cloud. We combine these measurements with molecular data observations from the ORION-B Large Program on the IRAM 30-meter telescope. We summarize our findings obtained using these observations as follows:
@@
We use HAWC+ Band D at µm to characterize the geometry of magnetic field across NGC 2024. We find that the structure of the magnetic field is ordered and follows the morphology of the expanding HII region and the direction of the filament.
@@
Using the CN(1–0) and HCO(1–0) molecular emission obtained using the IRAM 30-meter telescope, we characterize physical conditions (turbulence and gas density) in four specific regions in NGC 2024: the edges of the expanding HII shell located to the east and the west, the filament to the south and the environment in the northern part of NGC 2024. In our analysis, we include collisions with electrons and the hyperfine structure of CN(1–0), which are essential aspects of our calculations. Both CN(1–0) and HCO(1–0) emission lines are optically thick across NGC 2024, which broadens the observed line widths. The gas number density derived from CN(1–0) is comparable or somewhat higher than that obtained from analyzing the excitation of HCO(1–0).
@@
We derive the POS magnetic field strength using the ST (and classical DCF) method, with values ranging from 20 to 90 µG (50 - 240 µG) in the environments mentioned above. The strongest magnetic field is found in the region composed of the dusty filament and the edge of the expanding HII region, located in the northern part of NGC 2024. The high magnetic field strength derived in this area is driven by the largest gas densities and line widths.
@@
Magnetic field strengths derived using CN(1–0) and HCO(1–0) are comparable within the uncertainties at the edges of the bubble, especially in the eastern side of the bubble. Magnetic field strength calculated from HCO(1–0) is generally higher than that inferred from CN(1–0). measured from CN(1–0) and HCO(1–0) is larger on the eastern side of the bubble than on the western edge. This observed difference may result from changes in the magnetic field direction indicated in previous studies. In addition, the western side of the bubble is more impacted by stellar radiation due to its lower density.
@@
By analyzing mass-to-flux ratio, Alfvénic Mach number, and plasma-beta parameter, we find that the edges of the bubble and the filament show different properties. Gas impacted by the stellar feedback and traced by the CN(1–0) and HCO(1–0) emission seems better supported against gravitational collapse by the magnetic field than the gas in the filament, which represents a location where star formation can take place. However, we find the Alfvénic Mach number higher than 1 in all regions, which suggests that the magnetic field does not control the gas motions. Our results should be treated as upper limits given the use of the POS component of the -field. In addition, the estimated mass-to-flux ratio has large uncertainty due to the specific geometry of NGC 2024. We also note that these regions of NGC 2024 could be in the transition phase regarding gravitational stability.
@@ A.1 Filtering of low-level extended flux in HAWC+ images
Dust continuum images obtained with ground-based or airborne far-infrared total power continuum cameras suffer from some degree of correlated atmospheric noise, which cannot be completely separated from the extended low-level emission of the source. The HAWC+ instrument data reduction pipeline attempts to remove correlated components from the time stream using the algorithms described in Kovács (2008). To estimate the magnitude of the resulting spatial filtering at different flux levels in our HAWC+ Stokes images, we compared the Band D image to that obtained using the 160 µm channel of the PACS instrument on Herschel (OBSID 1342206080). Only high S/N ratio pixels are included in the dust polarization analysis (see Sec. 2.1). Fig. 6 shows a pixel-by-pixel comparison of the HAWC+ and PACS fluxes at the spatial resolution of the HAWC+ Band D data. Given the difference in the filter passbands and overall calibration uncertainties of the two instruments, we scaled the HAWC+ data by a factor of 1.12 to enforce a unity flux ratio at pixels with fluxes above 80,000 MJy sr (marked by the dotted line in the upper-right corner of the figure).
@@ B.1 Hyperfine structure of the CN line
Four parameters describe the hyperfine structure model: and . The first parameter is the antenna temperature multiplied by the optical depth of the CN emission:
@@
Before fitting, we define a two-dimensional mask based on the S/N ratio of the CN integrated intensity. We select the threshold of 10 and fit two CN components in the area that goes inside the mask and one component outside the mask. By adding this step, we provide CUBE with additional information about the area where the CN emission is bright enough to observe double components in its spectrum. CUBE requires the initial values of the free parameters for each component we want to fit. We specify our initial guess for the centroid velocities of the components to be 10 km/s and 7 km/s. However, we do not specify initial conditions for the rest of the fitting parameters and let CUBE find the best possible values. We show the results of the CN emission fitting for both components in Fig. 12.
@@ B.2 Gaussian lines of the HCO emission
The HCO emission line can be described using the Gaussian function. Similarly, as in the case of the CN, by inspecting the HCO data cube, we notice the presence of two velocity components across the large portion of the map. In addition, we observe a region in which HCO spectra contain extended line wings, suggestive of the molecular outflow also seen in the CO emission, whose presence is also known from previous studies (e.g., Richer et al. 1992). Different velocity components we observe correspond to the primary and intermediate velocity layers seen in the CO emission and its isotopologues across NGC 2024 (Gaudel et al. 2023).
@@ D.2 Histogram analysis
In this section, we describe an alternative approach to compute . We consider all points within each circularly-shaped region and create a histogram showing the distribution of magnetic field angles. Next, we compute the circular mean and fit the Gaussian function to the distribution of angles. We show the histograms for each beam-averaged area we used in this work, including the fitting parameters of each distribution in Fig. 24. In most cases, the histogram corresponds to a single Gaussian. In Fig. 25, we show comparison between from sliding window (y-axis) and histogram analysis (x-axis). The derived standard deviations from the sliding window (Sec. 5) are higher than those derived from the histogram analysis.

### 07 Feb 2024 v2

@@ 6.2.1 The edge of the HII region
In our work, we estimate the POS magnetic field strength at the border of the expanding HII region to be µG on the west side, and µG on the east side. The good agreement between the CN and HCO strengths in these regions, particularly at the eastern side, could indicate that these two molecular lines are tracing similar gas. Nevertheless, the factor of almost 2 difference in the measured in the west and the east could be due to different gas densities at these sides of the bubble. As the eastern side of the bubble is denser than the western side (see, Meyer et al., 2008), we expect that the gas here is less impacted by the incoming radiation. In addition, the opacity of the HCO emission is notably higher () on the western side than on the eastern side ().
@@ 6.2.2 Filament
Overall, our measurement of the magnetic field strength in the filament is lower than in other filamentary structures. For example, Pattle et al. (2017) measured a POS magnetic field strength of a few mG in the Orion A filament. Similarly, as discussed in the previous section, it is important to highlight that Pattle et al. (2017) used CO emission to measure the turbulence and number density and different method to derive the . Similarly, Ching et al. (2018) investigated magnetic field toward the DRL1 filament and found µG.
6.3 Magnetic field support in NGC 2024
@@
We report our results in Tab. 4 for values of magnetic field strengths computed from CN and HCO respectively using approach from Skalidis & Tassis (2021); Skalidis et al. (2021). We estimate in each region. This result could imply the presence of magnetically supported gas (Pattle et al., 2019). In general, it is worth noting that we see a difference of one and two orders of magnitude in the between the edges of the bubble and the filament and the overlap region.
@@
The results reported in Tab. 4 should be taken as upper limits considering we calculated them using the POS component of the magnetic field. In particular, the reported mass-to-flux ratio must be taken with caution. The complex geometry of NGC 2024, large uncertainties of reported POS magnetic field strength, and non availability of the LOS component of the -field have a big impact on values reported in Tab. 4. Our results imply that the edges of the bubble show different properties than the filamentary structure and the overlap region in NGC 2024. These regions are indeed different in terms of the impact of stellar feedback. However, further quantification of the gravitational stability of the gas in these regions requires systematic analysis of the magnetic field. This also includes both POS and LOS components in these regions and constraints on the geometry of NGC 2024, which is beyond the scope of this study.
The properties of the gas we study in our work vary with the environment and this gas shows different characteristics. Our results thus highlight the non-negligible role of the magnetic field in NGC 2024 in regions impacted by the stellar feedback. The edges of the bubble in NGC 2024 are magnetically dominated and gas in these regions is magnetically supported against gravitational collapse. However, the high values of the Alfvénic Mach number suggest that these regions are highly turbulent. The gravitational stability of gas located at the shell of expanding HII region is not always the case. For example, the expansion of the HII region can also trigger star formation, as seen in Galactic HII regions RCW 82 (Pomarès et al., 2009) and RCW 120 (Figueira et al., 2017). Since the HII region in NGC 2024 is relatively young, it is possible that the stellar feedback has not yet been strong enough to trigger star formation. Nevertheless, it is crucial to point out that our result is based on CN and HCO measurements, which trace the UV-illuminated and UV-shielded gas, which do not necessarily trace the star-forming gas.
@@ 7 Conclusions
We present new SOFIA HAWC+ dust polarization measurements across the NGC 2024 HII region and associated molecular cloud in the Orion B molecular cloud. We combined these measurements with molecular data observations from the ORION-B Large Program on the IRAM 30-meter telescope. We summarize our findings obtained using these observations as follows:
@@
We use HAWC+ Band D at µm to characterize the geometry of the magnetic field across NGC 2024. We find that the structure of the magnetic field is ordered and follows the morphology of the expanding HII region and the direction of the filament.
@@
Using the CN(1–0) and HCO(1–0) molecular emission obtained using the IRAM 30-meter telescope, we characterize physical conditions (turbulence and gas density) in four specific regions in NGC 2024: the edges of the expanding HII shell located to the east and the west, the filament to the south, and the environment in the northern part of NGC 2024. In our analysis, we include collisions with electrons and the hyperfine structure of CN(1–0), which are essential aspects of our calculations. Both CN(1–0) and HCO(1–0) emission lines are optically thick across NGC 2024, which broadens the observed line widths. The gas number density derived from CN(1–0) is comparable or somewhat higher than that obtained from analyzing the excitation of HCO(1–0).
@@
We derived the POS magnetic field strength using the ST (and classical DCF) method, with values ranging from 20 to 90 µG (50 - 240 µG) in the environments mentioned above. The strongest magnetic field is found in the region composed of the dusty filament and the edge of the expanding HII region, located in the northern part of NGC 2024. The high magnetic field strength derived in this area is driven by the largest gas densities and line widths.
@@
Magnetic field strengths derived using CN(1–0) and HCO(1–0) are comparable within the uncertainties at the edges of the bubble, especially in the eastern side of the bubble. The magnetic field strength calculated from HCO(1–0) is generally higher than that inferred from CN(1–0). We note that measured from CN(1–0) and HCO(1–0) is larger on the eastern side of the bubble than in the western edge. This observed difference may result from changes in the magnetic field direction indicated in previous studies. In addition, the western side of the bubble is more impacted by stellar radiation due to its lower density.
@@
By analyzing the mass-to-flux ratio, Alfvénic Mach number, and plasma-beta parameter, we find that the edges of the bubble and the filament show different properties. Gas impacted by the stellar feedback and traced by the CN(1–0) and HCO(1–0) emission seems better supported against gravitational collapse by the magnetic field than the gas in the filament, which represents a location where star formation can take place. However, we find the Alfvénic Mach number higher than one in all regions, which suggests that the magnetic field does not control the gas motions. Our results should be treated as upper limits given the use of the POS component of the field. In addition, the estimated mass-to-flux ratio has a large uncertainty due to the specific geometry of NGC 2024. We also note that these regions of NGC 2024 could be in the transition phase regarding gravitational stability.
@@ A.1 Filtering of low-level extended flux in HAWC+ images
Dust continuum images obtained with ground-based or airborne far-infrared total power continuum cameras suffer from some degree of correlated atmospheric noise, which cannot be completely separated from the extended low-level emission of the source. The HAWC+ instrument data reduction pipeline attempts to remove correlated components from the time stream using the algorithms described in Kovács (2008). To estimate the magnitude of the resulting spatial filtering at different flux levels in our HAWC+ Stokes images, we compared the Band D image to that obtained using the 160 µm channel of the PACS instrument on Herschel (OBSID 1342206080). Only high S/N pixels are included in the dust polarization analysis (see Sec. 2.1). Fig. 6 shows a pixel-by-pixel comparison of the HAWC+ and PACS fluxes at the spatial resolution of the HAWC+ Band D data. Given the difference in the filter passbands and overall calibration uncertainties of the two instruments, we scaled the HAWC+ data by a factor of 1.12 to enforce a unity flux ratio at pixels with fluxes above 80,000 MJy sr (marked by the dotted line in the upper-right corner of the figure).
@@ B.1 Hyperfine structure of the CN line
Four parameters describe the hyperfine structure model: , and . The first parameter is the antenna temperature multiplied by the optical depth of the CN emission:
@@
Before fitting, we define a two-dimensional mask based on the S/N of the CN integrated intensity. We select the threshold of 10 and fit two CN components in the area that goes inside the mask and one component outside the mask. By adding this step, we provide CUBE with additional information about the area where the CN emission is bright enough to observe double components in its spectrum. CUBE requires the initial values of the free parameters for each component we want to fit. We specify our initial guess for the centroid velocities of the components to be 10 km/s and 7 km/s. However, we do not specify initial conditions for the rest of the fitting parameters and let CUBE find the best possible values. We show the results of the CN emission fitting for both components in Fig. 12.
@@ B.2 Gaussian lines of the HCO emission
The HCO emission line can be described using the Gaussian function. Similarly, as in the case of the CN, by inspecting the HCO data cube, we notice the presence of two velocity components across the large portion of the map. In addition, we observe a region in which HCO spectra contain extended line wings, suggestive of the molecular outflow also seen in the CO emission, whose presence is also known from previous studies (for instance, see, Richer et al. 1992). Different velocity components we observe correspond to the primary and intermediate velocity layers seen in the CO emission and its isotopologues across NGC 2024 (Gaudel et al. 2023).
@@ D.2 Histogram analysis
In this section, we describe an alternative approach to compute . We consider all points within each circularly shaped region and create a histogram showing the distribution of magnetic field angles. Next, we compute the circular mean and fit the Gaussian function to the distribution of angles. We show the histograms for each beam-averaged area we used in this work, including the fitting parameters of each distribution in Fig. 24. In most cases, the histogram corresponds to a single Gaussian. In Fig. 25, we show comparison between from sliding window (y-axis) and histogram analysis (x-axis). The derived standard deviations from the sliding window (Sec. 5) are higher than those derived from the histogram analysis.